OpenStudy (countrygirl1431):

Simplify 5(x - 2) + 3x 8x - 2 8x - 10 -7x

OpenStudy (anonymous):

Hello there. So for this, you need to use the distributive property. do you remember what that is?

OpenStudy (countrygirl1431):

yeah

OpenStudy (anonymous):

Great so do the distributive property on 5(x-2) Ignore the +3x, we can add that in later.

OpenStudy (countrygirl1431):

so 5x-2

OpenStudy (countrygirl1431):

and then you would subrtact 3x from both sides

OpenStudy (anonymous):

Almost. Remember that a*(b*c) = (a*b) + (a*c) So for this question you did the first part...

OpenStudy (countrygirl1431):

8x-2

OpenStudy (anonymous):

You also need to distribute the 5 to the 2.

OpenStudy (anonymous):

5(x-2) So 5 is the a x is the b and -2 is the c (5*x) + (5 * -2)

OpenStudy (anonymous):

What you should get is (5x) + (-10)

OpenStudy (anonymous):

That's basically 5x - 10. Are you following so far?

OpenStudy (countrygirl1431):

yes

OpenStudy (anonymous):

Cool so now all we have left is the +3x which was not part of the distribution. So 5x + 3x = 8x. Our answer will be 8x - 10. Don't forget to distribute the second part.

OpenStudy (countrygirl1431):

ok thanks so much

OpenStudy (anonymous):

Yup. good luck!